OpenStudy (anonymous):

The function Q(t)=Qsub(o)e^-kt may be used to model radioactive decay. Q represents the quantity remaining after t years; k is the decay constant. What is the decay constant for plutonium-240 if its half-life is 6,300 years? A.-[ln(0.5)/6,300] B.ln(0.5)/ln(6,300) C.240ln(0.5)/6,300 D.ln(0.5)/240

OpenStudy (anonymous):

A is my reasoning correct lol or even the answer ????? e^(-6300k) = 0.5 -6300k = ln(0.5) k = ln(0.5) / (-6300) k = -ln(0.5) / 6300

Miracrown (miracrown):

yes,  So it looks like A is our winner! :)